# Sequential Monte Carlo Squared for Online Inference in Stochastic Epidemic Models

### v1

@@ Abstract
Effective epidemic modeling and surveillance require computationally efficient methods that can continuously update estimates as new data becomes available. This paper explores the application of an online variant of Sequential Monte Carlo Squared (O-SMC2) to the stochastic Susceptible-Exposed-Infectious-Removed (SEIR) model for real-time epidemic tracking. The particularity of O-SMC2 lies in its ability to update the parameters using a particle Metropolis-Hastings kernel, ensuring that the target distribution remains invariant while only utilizing a fixed window of recent observations. This feature enables timely parameter updates and significantly enhances computational efficiency compared to the standard SMC2, which processes the entire dataset. First, we demonstrate the efficiency of O-SMC2 on simulated data, where both the parameters and the observation process are known. We then apply the method to a real-world COVID-19 dataset from Ireland, successfully tracking the epidemic trajectory and estimating the time-dependent reproduction number of the disease. Our results show that O-SMC2 provides highly accurate online estimates of both static and dynamic epidemiological parameters while substantially reducing computational costs. These findings highlight the potential of O-SMC2 for real-time epidemic monitoring and supporting adaptive public health interventions.
@@ 1 Introduction
Disease modeling plays a crucial role in public health by providing insights into the mechanisms behind the spread of infectious diseases. Such understanding is vital for designing effective intervention strategies and supporting evidence-based policy-making. As the global landscape of infectious diseases evolves and the risk of novel outbreaks increases (Marani et al., 2021), there is a pressing need for robust modeling frameworks that can adapt to changing data, update parameter values as new information becomes available, and deliver accurate assessments and predictions of epidemic trajectories (Birrell et al., 2017).
Bayesian approaches have emerged as powerful tools for tackling these challenges, particularly in sampling from the posterior distribution of model parameters. Markov chain Monte Carlo (MCMC) methods have been widely utilized in epidemic analysis (O’Neill and Roberts, 1999; Jewell et al., 2009; Wang and Walker, 2022), enabling exact inference but requiring an explicit expression for the complete-data likelihood. Furthermore, these methods can be computationally intensive and inefficient for real-time applications, as they necessitate reanalyzing the entire dataset with each new data point (Birrell et al., 2017). An alternative approach is Approximate Bayesian Computation (ABC), which has gained popularity for analyzing epidemic processes due to its ability to bypass the need for explicit likelihood formulations (Kypraios et al., 2017). While ABC offers flexibility, it relies on summary statistics, which can introduce bias and limit its precision (Kypraios et al., 2017). Furthermore, ABC does not inherently address the challenge of sequentially updating estimates as new data are collected.
@@
Sequential Monte Carlo (SMC) methods, also known as particle filtering, offer an attractive solution to update the posterior distribution of a stochastic process as new data become available (Doucet et al., 2001). These methods have seen increasing application in epidemiology, with early studies demonstrating their effectiveness in calibrating nonlinear compartmental infectious disease models (Yang et al., 2014; Camacho et al., 2015). For instance, Welding and Neal (2019) employed a sequential Bayesian framework to the 2001 UK Foot-and-Mouth epidemic, utilizing MCMC algorithms to rejuvenate particles and facilitate real-time outbreak analysis through data augmentation. Additionally, Birrell et al. (2020) proposed an age-structured SEIR model for influenza outbreaks, showcasing the advantages of SMC over MCMC for online prediction, particularly in integrating multiple data sources for parameter inference.
@@
However, it is important to recognize that traditional SMC algorithms are not well suited for estimating static parameters (Kantas et al., 2015). Although extensions like particle MCMC (PMCMC) (Andrieu et al., 2010) have been developed to achieve exact sampling from the posterior by creating unbiased estimators for the likelihood through simulation and subsequently applying traditional MCMC methods, they remain computationally demanding. Each MCMC iteration typically requires a full SMC run, making these methods less suitable for real-time applications (Dureau et al., 2013; Funk et al., 2018; Endo et al., 2019; Cazelles et al., 2021). In their recent study, Storvik et al. (2023) propose an SMC framework for inferring a time-varying reproduction number of COVID-19 in Norway. It is essential to highlight that this SMC algorithm relies on a set of sufficient statistics to sequentially estimate hyperparameters that control the evolution of the reproduction number. However, the estimation of the posterior parameter relies on the assumption that sufficient statistics-based filters are analytically tractable, which will only hold for particular dynamical system setups.
@@
A promising solution to these challenges is the SMC2 method (Chopin et al., 2013; Jacob, 2015), which enhances traditional particle filtering by integrating a nested filtering mechanism. This approach simultaneously estimates model states and parameters, improving both the accuracy and adaptability of the estimation process. A key feature of the SMC2 algorithm is the “rejuvenation” step, which uses a PMCMC kernel to mitigate particle impoverishment. However, this step traditionally requires computing the likelihood using all past data points, which may limit its online applicability.
In this paper, we implement an online variant of the SMC2 method (O-SMC2), which reduces computational costs per iteration by targeting an approximate distribution and evaluating the PMCMC over a fixed window of recent observations (Vieira, 2018). This adaptation preserves the sequential and online attributes of the method, emphasizing its practical application in real-time state and parameter estimation for epidemic models, specifically the stochastic SEIR model.
@@
The remainder of this paper is structured as follows. Section 2 describes the state-space framework for the SEIR compartmental model and provides a brief overview of the SMC2/O-SMC2 methodology. In Section 3, we apply the O-SMC2 algorithm to both simulated data and real-world COVID-19 data from Ireland, demonstrating its potential for real-time epidemiological surveillance and for inferring the time-dependent reproduction number. Finally, Section 4 concludes the paper by summarizing the main findings and discussing the potential challenges and limitations associated with our model.
@@ 2.1 Stochastic SEIR model
Compartmental models for infectious diseases, which had their conceptual origins in the early to mid $20^{th}$ century, attempt to describe a disease’s progression and spread through individuals moving through disease stages over time. The stochastic SEIR model serves as the foundation for our investigation. Consider the time interval $(t,t+\delta t]$, where $\delta t$ represents the time between observations. If we assume that the time spent by an individual in a compartment is exponentially distributed with a rate $\lambda$, then the number of individuals leaving a compartment during the time step $\delta t$ can be seen as a Binomial variable with the probability $p=1-e^{-\lambda\delta t}$. This parameterization ensures that the probability will remain between 00 and $1$. Hence the discrete-time SEIR stochastic model is given by:
$S_{t+\delta t}$ $=S_{t}-Y_{SE}(t),$ $Y_{SE}(t)$ $\sim\mathrm{Bin}\left(S_{t},1-e^{-\beta_{t}\frac{I_{t}}{N}\delta t}\right)$
@@
Here, $S_{t}$, $E_{t}$, $I_{t}$, and $R_{t}$ represent the compartments of susceptible, exposed (infected but not yet infectious), infected, and removed individuals, respectively, at time $t$. Susceptible individuals, upon contact with infected individuals, transition to the exposed class at a rate $\beta_{t}$ (which can be time-dependent or constant). Exposed individuals become infectious at a rate $\sigma$, and infected individuals recover from the disease at a rate $\gamma$. The population is assumed to have a constant size $N=S_{t}+E_{t}+I_{t}+R_{t}$ at each time point.
Figure 1: Graphical representation of a SEIR model. Each arrow expresses the rate of transition between the compartments. $\beta_{t}$ is the disease transmission rate, $1/\sigma$ is the mean incubation period, and $1/\gamma$ is the mean infectious period.
@@ 2.2 State-space framework
In the context of the SEIR model, the system can be formulated as a state-space model (SSM), which consists of two stochastic processes: the first process governs the evolution of the disease within the population, while the second process represents the observation process. The state-space model is defined for time steps $t=1,\ldots,T$ as follows:
[…]
Here, $x_{t}$ denotes the latent state variables (state variable in the classical SEIR model), which evolve according to a Markov process with transition distribution $f(x_{t}|x_{t-1},\theta)$ given by the model (2.1). The observation $y_{t}$ is assumed to be probabilistically related to the hidden state $x_{t}$ via the distribution $g(y_{t}|x_{t},\theta)$. The parameter vector $\theta$ encompasses all static epidemiological parameters of the model such as $\beta$, $\sigma$ and $\gamma$, governing both the state transitions and the observation process. Throughout the study, $\theta$ will be used interchangeably to refer to these parameters.
@@ 2.3 State and Parameter inference
In this section, we briefly describe the procedure for sequential parameter estimation using the SMC2 scheme. Let $x_{0:t}$  [^1] For any sequence $\{z_{t}\}_{t\geq 0}$, we denote $z_{i:j}=\{z_{i},z_{i+1},\ldots,z_{j}\}$ and $z^{1:N}_{j}$ (resp. $z^{1:N}_{i:j}$) is a set of $N$ random variables $z^{l}_{j}$ (resp. $z^{l}_{i:j}$), for $l=1,...,N$. represent the sequence of latent states (in our case the SEIR compartment counts) up to time $t$, and let $y_{1:t}$ represent the observations (e.g., daily reported cases). The goal is to construct a sequence of posterior distributions $\{p(\theta|y_{1:t}),t=1,\ldots,T\}$ that reflect the evolution of parameter estimates as new observations become available. Bayes’ rule provides the basis for updating the posterior distribution as follows:
@@
which suggests a sequential importance sampling approach. This approach involves reweighting a set of parameter particles based on the incremental likelihood $p(y_{t}|y_{1:t-1},\theta)$. The iterated batch importance sampling (IBIS) method (Chopin, 2002) employs this technique with an MCMC step to rejuvenate parameter samples. However, for state-space models, the incremental likelihood is generally intractable. To address this issue, we can replace $p(y_{t}|y_{1:t-1},\theta)$ with an unbiased estimate derived from running a particle filter, targeting the filtering distribution $p(x_{t}|y_{1:t},\theta)$ for each of the $\theta$-particles. This approach is known as SMC2 due to the use of nested filters. For a comprehensive discussion on this approach, including formal justifications and recent developments, refer to (Chopin et al., 2013; Jacob, 2015).
@@ 2.3.1 Particle filtering
For a given value of the model parameters $\theta$, we aim to infer the latent variables $x_{t}$ at each time point, given the observations $y_{1:t}$ by means of the filtering distribution $p(x_{t}|y_{1:t},\theta)$. Particle filtering is a technique that approximates this filtering distribution by sequentially generating a set of particles with associated weights $\{x_{t}^{i},w_{t}^{i}\}_{i=1}^{N_{x}}$, via importance sampling and resampling. Hence, for a known parameter value $\theta$, the posterior distribution of the latent states is represented as a Dirac delta mixture:
@@
One notable particle filtering method is the Bootstrap particle filter (BPF) (Gordon et al., 1993), which simplifies weight calculations by setting the importance distribution equal to the state transition distribution, i.e, $q(x_{t}|x_{t-1},y_{t},\theta)=f(x_{t}|x_{t-1},\theta)$. Other advanced filtering algorithms have been developed to improve efficiency (Pitt and Shephard, 1999), but these are not discussed in this study. The general procedure for the Bootstrap Particle Filter is shown in Algorithm 1.
@@
Output: Particles set :$\left\{x^{i}_{0:t},w^{i}_{0:t}\right\}_{i=1}^{N_{x}}$
@@
5:     Propagate states $x_{t}^{i}\sim f(\cdot|x_{t-1}^{a_{t}^{i}},\theta)$
@@
The resampling step in Algorithm 1, referred to in step 4, aims to mitigate particle degeneracy [^2] Particle degeneracy refers to the situation where, after many iterations, most particle weights approach zero while only a few retain significant weight. by discarding particles with lower weights and duplicating those with higher weights. After resampling, we reset the weights to $w^{i}_{t-1}=1$. Various resampling schemes exist as described in (Doucet et al., 2009). Unless stated otherwise, we use stratified resampling in all our implementations due to its efficiency and lower computational cost (Sheinson et al., 2014; Vieira, 2018). In Algorithm 1, if the resampling step is omitted, then the weights are replaced by $w^{i}_{t}=w^{i}_{t-1}g(y_{t}|x_{t}^{i},\theta)$. Particle filtering can also be used for Bayesian parameter inference by providing an unbiased estimate of the incremental likelihood:
@@
Hence, by the law of total probability, we can deduce an SMC estimate of the marginal likelihood for a given parameter $\theta$:

### v2

@@ Abstract
Effective epidemic modeling and surveillance require computationally efficient methods that can continuously update parameter estimates as new data becomes available. This paper explores the application of an online variant of Sequential Monte Carlo Squared (O-SMC2) to the stochastic Susceptible-Exposed-Infectious-Removed (SEIR) model for real-time epidemic tracking. The advantage of O-SMC2 lies in its ability to update parameter estimates using a particle Metropolis-Hastings kernel by only utilizing a fixed window of recent observations. This feature enables timely parameter updates and significantly enhances computational efficiency compared to standard SMC2, which by comparison, requires processing of all past observations. First, we demonstrate the efficiency of O-SMC2 on simulated epidemic data, where both the true parameter values and the observation process are known. We then make an application to a real-world COVID-19 dataset from Ireland, successfully tracking the epidemic and estimating a time-dependent reproduction number of the disease. Our results show that O-SMC2 provides highly accurate online estimates of both static and dynamic epidemiological parameters while substantially reducing computational cost. These findings highlight the potential of O-SMC2 for real-time epidemic monitoring and supporting adaptive public health interventions.
@@ 1 Introduction
Disease modeling plays a crucial role in public health by providing insights into the mechanisms behind the spread of infectious diseases. Such understanding is vital for designing effective intervention strategies and supporting evidence-based policy-making. As the global landscape of infectious diseases evolves and the risk of novel a pandemic events like COVID-19 increases (Marani et al., 2021), there is a pressing need for robust modeling frameworks that can adapt to changes in the underlying disease dynamics, progressively update parameter values as new information becomes available, and deliver accurate assessments and predictions of epidemic trajectories (Birrell et al., 2017).
Traditionally, parameter inference in epidemic models has been approached using both frequentist and Bayesian methods. Frequentist techniques such as maximum likelihood estimation (Althaus, 2014) and nonlinear least squares (Aloke et al., 2023) have been widely applied to deterministic models based on ordinary differential equations due to their simplicity and computational efficiency. Bayesian approaches, have gained popularity for their ability to incorporate prior knowledge and quantify uncertainty. In this context, Markov chain Monte Carlo (MCMC) methods remain a popular choice for sampling from posterior distributions of model parameters (O’Neill and Roberts, 1999; Jewell et al., 2009; Wang and Walker, 2022). Recent advances in probabilistic programming tools like Stan (Carpenter et al., 2017; Grinsztajn et al., 2021) and specialized toolboxes such as quantdiffforecast (Chowell et al., 2024) and Bayesianfitforecast (Karami et al., 2024) have made Bayesian workflows more accessible and flexible in epidemiological modeling. However, the computational demands of MCMC limit its utility for real-time applications. This is because incorporating incrementally observed new data points typically necessitates re-running the entire inference process (Birrell et al., 2017). When the likelihood function is intractable but simulation from the generative model is feasible, likelihood-free approaches such as Approximate Bayesian Computation (ABC) are commonly used (Kypraios et al., 2017). While ABC methods offer flexibility and an intuitive framework, their performance relies heavily on a user chosen set of summary statistics to measure the distance between simulated and observed data. Poorly chosen, these can detrimentally impact performance. Standard ABC lacks native mechanisms for sequential parameter updating as new data arrive incrementally (Minter and Retkute, 2019). Sequential adaptations of ABC are possible (Toni et al., 2009), however these still depend heavily on user chosen summary statistics, creating challenges in complex modelling scenarios (Han et al., 2025).
@@
We adopt an approach which gets around many of the limitations outlined above, and enables sequential and probabilistically well calibrated inference. Sequential Monte Carlo (SMC) methods, also known as particle filtering, provide a natural framework for real-time state and parameter estimation by sequentially updating a posterior distribution as new data becomes available (Doucet et al., 2001). These methods have seen growing application in epidemiology, with early studies demonstrating their effectiveness in calibrating nonlinear compartmental models of infectious disease dynamics (Yang et al., 2014; Camacho et al., 2015). For example, Welding and Neal (2019) applied a sequential Bayesian framework to the 2001 UK Foot-and-Mouth epidemic, using MCMC-based rejuvenation steps to mitigate particle degeneracy and enable real-time outbreak analysis. Similarly, Birrell et al. (2020) developed an age-structured SEIR model on simulated influenza outbreak data, demonstrating efficient real-time inference by integrating multiple data sources and reducing computational costs compared to standard MCMC.
@@
While SMC methods are well-suited for sequential state estimation, they are not directly able to facilitate the estimation of static parameters (Kantas et al., 2015). Hybrid techniques such as particle MCMC (PMCMC) (Andrieu et al., 2010) address this limitation by combining SMC with MCMC. PMCMC generates unbiased likelihood estimates through particle filtering, enabling exact posterior sampling via traditional MCMC methods. However, these methods remain computationally demanding since each MCMC iteration requires a full SMC run, limiting their suitability for real-time applications (Dureau et al., 2013; Funk et al., 2018; Endo et al., 2019; Cazelles et al., 2021). In their recent study, Storvik et al. (2023) propose an SMC framework for inferring a time-varying reproduction number of COVID-19 in Norway. Their approach sequentially estimates hyperparameters controlling the reproduction number by relying on a set of sufficient statistics. However, this approach assumes that the sufficient statistics-based filters are analytically tractable, a condition that only holds for particular types of dynamical systems.
@@
A promising alternative is the SMC2 algorithm (Chopin et al., 2013; Jacob, 2015; Rosato et al., 2023), which integrates nested particle filters to simultaneously estimate model states and parameters, improving both accuracy and adaptability. A key step in SMC2 is the so-called “rejuvenation” using PMCMC kernels to prevent particle impoverishment. Nevertheless, this step conventionally requires computing the likelihood over all past observations, posing challenges for truly online or streaming data applications.
To address this limitation, Vieira (2018) introduced a computationally efficient modification that evaluates the PMCMC over a fixed-size moving window of recent observations. In this study, we apply and evaluate this fixed-window SMC2 approach, referred to as online SMC2 (O-SMC2), within the context of epidemic modeling. While O-SMC2 is an established algorithm, its strengths for disease modeling have not been explored. Our contribution is its tailored implementation and comprehensive assessment for sequential state and parameter inference in stochastic SEIR models under real-time constraints. This is an area of growing practical importance for epidemic response planning. Compared to standard SMC2, the O-SMC2 framework substantially reduces computational demands while preserving the capacity for accurate sequential estimation. This efficiency gain is particularly relevant for fast-moving outbreaks, where timely estimates of key epidemiological parameters such as the time-varying effective reproduction number and the infectious period are critical for guiding public health interventions.
@@
The remainder of this paper is structured as follows. Section 2 describes the state-space framework for the SEIR compartmental model, provides a brief overview of the SMC2/O-SMC2 methodology, and discusses parameter identifiability. In Section 3, we apply the O-SMC2 algorithm to both simulated data and real-world COVID-19 data from Ireland, demonstrating its potential for real-time epidemiological surveillance and inference of the time-dependent reproduction number. Finally, Section 4 concludes the paper by summarizing the main findings and discussing potential challenges and limitations of our approach.
@@ 2.1 Stochastic SEIR model
Compartmental models for infectious diseases, which had their conceptual origins in the early to mid $20^{th}$ century, attempt to describe a disease’s progression and spread through individuals moving through disease stages over time. The stochastic SEIR model serves as the foundation for our investigation. Consider the time interval $(t,t+\delta t]$, where $\delta t$ represents the time between observations. If we assume that the time spent by an individual in a compartment is exponentially distributed with a rate $\lambda$, then the number of individuals leaving a compartment during the time step $\delta t$ can be seen as a Binomial random variable with success probability $p=1-e^{-\lambda\delta t}$ (this parameterization ensures a success probability between 0 and $1$, see Lekone and Finkenstädt (2006)). The discrete-time SEIR stochastic model is specified by:
$S_{t+\delta t}$ $=S_{t}-Y_{SE}(t),$ $Y_{SE}(t)$ $\sim\mathrm{Bin}\left(S_{t},1-e^{-\beta\frac{I_{t}}{N}\delta t}\right)$
@@
Here, $S_{t}$, $E_{t}$, $I_{t}$, and $R_{t}$ represent the compartments of susceptible, exposed (infected but not yet infectious), infectious, and removed individuals, respectively, at time $t$. Susceptible individuals, upon contact with infectious individuals, transition to the exposed class at a rate $\beta$ (which can be time-dependent or constant). Exposed individuals become infectious at a rate $\sigma$, and infectious individuals recover from the disease at a rate $\gamma$ (Figure 1). The population is assumed to have a constant size $N=S_{t}+E_{t}+I_{t}+R_{t}$ at each time point.
Figure 1: Graphical representation of a SEIR model. Each arrow expresses the rate of transition between the compartments. $\beta$ is the disease transmission rate, $1/\sigma$ is the mean incubation period, and $1/\gamma$ is the mean infectious period.
@@ 2.2 State-space framework
In the context of the SEIR model, the system can be formulated as a state-space model (SSM), which consists of two stochastic processes: one governing the evolution of the disease within the population, and the other describing the observation process (Figure 2). The state-space model is defined for time steps $t=1,\ldots,T$ as follows:
$x_{0}\sim f(x_{0}|\theta),$ $\rhd$ Initial state (2)
$x_{t}|x_{t-1}\sim f(x_{t}|x_{t-1},\theta),$ $\rhd$ State transition process (3)
$y_{t}|x_{t}\sim g(y_{t}|x_{t},\theta).$ $\rhd$ Observation process (4)
Here, $x_{t}$ denotes the vector of latent state variables (state variable in the classical SEIR model), which evolve according to a Markov process with transition distribution $f(x_{t}|x_{t-1},\theta)$ given by the model (2.1). The observation $y_{t}$ is assumed to be probabilistically related to the hidden state $x_{t}$ via the distribution $g(y_{t}|x_{t},\theta)$. The parameter vector $\theta$ encompasses all static epidemiological parameters of the model such as $\beta$, $\sigma$ and $\gamma$, governing both the state transitions and the observation process. Throughout the study, $\theta$ refers collectively to all state space model and observation process parameters.
@@ 2.3 State and Parameter inference
In this section, we briefly describe the procedure for sequential parameter estimation using the SMC2 scheme. Let $x_{0:t}$  [^1] For any sequence $\{y_{t}\}_{t\geq 0}$, we denote $z_{i:j}=\{z_{i},z_{i+1},\ldots,z_{j}\}$ and $z^{1:N}_{j}$ (resp. $z^{1:N}_{i:j}$) is a set of $N$ random variables $z^{l}_{j}$ (resp. $z^{l}_{i:j}$), for l=1,,Nl=1,...,N. represent the sequence of latent states (in our case the SEIR compartment counts) up to time $t$, and let $y_{1:t}$ represent the observations (e.g., daily reported cases). The goal is to construct a sequence of posterior distributions $\{p(\theta|y_{1:t}),t=1,\ldots,T\}$ that reflect the evolution of parameter estimates as new observations become available. Bayes’ rule provides the basis for updating the posterior distribution as follows:
@@
which suggests a sequential importance sampling approach. This approach involves reweighting a set of parameter particles based on the incremental likelihood $p(y_{t}|y_{1:t-1},\theta)$. The iterated batch importance sampling (IBIS) method (Chopin, 2002) employs this technique with an MCMC step to rejuvenate parameter samples. However, for state-space models, the incremental likelihood is generally intractable. To address this issue, we can replace $p(y_{t}|y_{1:t-1},\theta)$ with an unbiased estimate derived from running a particle filter, targeting the filtering distribution $p(x_{t}|y_{1:t},\theta)$ for each of the $\theta$-particles. This approach is known as SMC2 due to the use of nested filters. For a comprehensive discussion on this approach, including formal justifications, refer to (Chopin et al., 2013).
@@ 2.3.1 Particle filtering
For a given value of the model parameters $\theta$, the latent variables $x_{t}$ at each time point, can be inferred from the observations $y_{1:t}$ using the filtering distribution $p(x_{t}|y_{1:t},\theta)$. Particle filtering is a technique that approximates this filtering distribution by sequentially generating a set of particles with associated weights $\{x_{t}^{i},w_{t}^{i}\}_{i=1}^{N_{x}}$, via importance sampling and resampling. Hence, for a known parameter value $\theta$, the posterior distribution of the latent states is represented as a Dirac delta mixture:
@@
One notable particle filtering method is the Bootstrap particle filter (BPF) (Gordon et al., 1993), which simplifies weight calculations by setting the importance distribution equal to the state transition distribution, i.e, $q(x_{t}|x_{t-1},y_{t},\theta)=f(x_{t}|x_{t-1},\theta)$. Many filtering algorithms have been developed to improve efficiency (Pitt and Shephard, 1999), but for this study, we focus on the BPF. The general procedure for the Bootstrap Particle Filter is shown in Algorithm 1.
@@
Output: Particle set $\left\{x^{i}_{0:t},w^{i}_{0:t}\right\}_{i=1}^{N_{x}}$
@@
5:  Propagate states $x_{t}^{i}\sim f(|x_{t-1}^{a_{t}^{i}},\theta)$
@@
The resampling step in Algorithm 1, referred to in step 4, aims to mitigate particle degeneracy [^2] Particle degeneracy refers to the situation where, after many iterations, only a small number of particles dominate and retain a large proportion of the weight. by discarding particles with lower weights and duplicating those with higher weights. After resampling, we reset all weights to $w^{i}_{t-1}=1$. Various resampling schemes exist as described in Doucet et al. (2009). Unless stated otherwise, we use stratified resampling in all our implementations due to its efficiency and lower computational cost (Sheinson et al., 2014; Vieira, 2018). In Algorithm 1, if the resampling step is omitted, then the weights are replaced by $w^{i}_{t}=w^{i}_{t-1}g(y_{t}|x_{t}^{i},\theta)$. Particle filtering can also be used for Bayesian parameter inference since it provides an unbiased estimate of the incremental likelihood:
@@
Using the law of total probability, an SMC estimate of the marginal likelihood at $\theta$ follows: